# Periodic nonlinear Schrödinger equation with distributional potential and invariant measures

### v1

@@ Abstract
$\sigma_{\kappa}(p)=\frac{1}{3}-\frac{\kappa}{6}$ (3.13)
when $p\leq 8$ and
$\sigma_{\kappa}(p)=\frac{4-\kappa}{6}-\frac{2}{p-2}$ (3.14)
when $p>8$.
@@
Let $V\in\mathcal{C}^{-1+\kappa}$ and $p\geq 2$. For any $\sigma>\sigma_{\kappa}(p)$, there exists a unique solution $u\in C([0,T],H^{\sigma})$ to equation (1.1) for $u_{0}\in H^{\sigma}$ up to time $T$ of order $\|u_{0}\|_{H^{\sigma}}^{-\frac{6(p-2)}{p-8}}$ and we have
$\sup_{t\in[0,T]}\|u(t)\|_{H^{\sigma}}\leq 2\|u_{0}\|_{H^{\sigma}}.$ (3.15)
Moreover, we have
$\sup_{t\in[0,T]}\|u(t)-v(t)\|_{H^{\sigma}}\lesssim\|u_{0}-v_{0}\|_{H^{\sigma}}$ (3.16)
for two solutions $u$ and $v$ with respective initial data $u_{0}$ and $v_{0}$ with $T$ of order $\|u_{0}\|_{H^{\sigma}}^{-\frac{6(p-2)}{p-8}}\wedge\|v_{0}\|_{H^{\sigma}}^{-\frac{6(p-2)}{p-8}}$.
@@
$i\partial_{t}v=\mathcal{H}^{\sharp}v+\lambda\Gamma^{-1}(|\Gamma v|^{p-2}\Gamma v)$ (3.17)
with initial data $v(0)=\Gamma^{-1}u_{0}\in H^{\sigma}$. We prove existence and uniqueness of solutions for this equation, this will gives a unique solution to the initial equation using that $\Gamma$ is continuous from $H^{\sigma}$ to itself for $\sigma\leq 1+\kappa$. For $T>0$, consider the solution space
@@
for $\sigma\in(\sigma_{\kappa},\frac{1}{2})$ and $\varepsilon\in(0,\sigma-\sigma_{\kappa}]$. We prove that the map
$\Phi(v)(t):=e^{-it\mathcal{H}^{\sharp}}v_{0}+i\lambda\int_{0}^{t}e^{-i(t-s)\mathcal{H}^{\sharp}}\Gamma^{-1}\big{(}|\Gamma v(s)|^{p-2}\Gamma v(s)\big{)}\mathrm{d}s$ (3.20)
@@
$\|\Phi(v)(t)\|_{H^{\sigma}}$ $\lesssim\|v_{0}\|_{H^{\sigma}}+\int_{0}^{t}\||\Gamma v(s)|^{p-2}\Gamma v(s)\|_{H^{\sigma}}\mathrm{d}s$
$\lesssim\|v_{0}\|_{H^{\sigma}}+\int_{0}^{t}\|v(s)\|_{H^{\sigma}}\|v(s)\|_{L^{\infty}}^{p-2}\mathrm{d}s$
$\lesssim\|v_{0}\|_{H^{\sigma}}+\|v\|_{L^{\infty}([0,t],H^{\sigma})}T^{1-\frac{p-2}{6}}\|v\|_{L^{6}([0,t],L^{\infty})}^{p-2}$
using that $\Gamma$ is continuous from $H^{\sigma}$ to itself, from $L^{\infty}$ to itself and Hölder inequality for $p<8$ for $t\in[0,T]$. Since $\sigma<\frac{1}{2}$, the space $H^{\sigma}$ is not an algebra and one needs to control the $L^{\infty}$ norm, this is where Strichartz inequalities are crucial with the embedding
@@
$\sup_{t\in[0,T]}\|\Phi(v)(t)\|_{H^{\sigma}}\lesssim\|v_{0}\|_{H^{\sigma}}+T^{\frac{8-p}{6}}\|v\|_{\mathcal{S}_{T}}^{p-1}.$ (3.22)
Since the spaces $\mathcal{D}^{\delta}$ and $H^{\delta}$ are equivalent for $|\delta|\leq 1+\kappa$ and using the continuity results on $\Gamma$, we have
@@
$\|\Phi(v)\|_{L^{6}([0,t],W^{\frac{1}{6}+\varepsilon,6})}$ $\lesssim\|v_{0}\|_{H^{\frac{1-\kappa}{6}+\varepsilon+\frac{1}{6}}}+\int_{0}^{t}\|v(s)\|_{H^{\frac{1-\kappa}{6}+\varepsilon+\frac{1}{6}}}\|v(s)\|_{L^{\infty}}^{p-2}\mathrm{d}s$ (3.24)
$\lesssim\|v_{0}\|_{H^{\frac{1-\kappa}{6}+\varepsilon+\frac{1}{6}}}+\|v\|_{L^{\infty}([0,t],H^{\frac{1-\kappa}{6}+\varepsilon+\frac{1}{6}})}T^{1-\frac{p-2}{6}}\|v\|_{L^{6}([0,t],W^{\frac{1}{6}+\varepsilon,6})}^{p-2}$ (3.25)
$\lesssim\|v_{0}\|_{H^{\sigma}}+T^{\frac{8-p}{6}}\|v\|_{\mathcal{S}_{T}}^{p-1}$ (3.26)
since $\sigma\geq\sigma_{\kappa}+\varepsilon$. This gives
$\|\Phi(v)\|_{\mathcal{S}_{T}}\lesssim\|v_{0}\|_{H^{\sigma}}+T^{\frac{8-p}{6}}\|v\|_{\mathcal{S}_{T}}^{p-1}$ (3.27)
hence for any $A\geq 0$, the ball centered in 00 of size $A$ is stable by $\Phi$ for $T>0$ small enough. The same computation gives
$\|\Phi(v)-\Phi(v^{\prime})\|_{\mathcal{S}_{T}}\lesssim T^{\frac{8-p}{6}}(1+\|v\|_{\mathcal{S}_{T}}^{p-2}+\|v^{\prime}\|_{\mathcal{S}_{T}}^{p-2})\|v-v^{\prime}\|_{\mathcal{S}_{T}}$ (3.28)
for $v,v^{\prime}\in\mathcal{S}_{T}$ with $v(0)=v^{\prime}(0)$. This gives that for any $A>0$ and $v_{0}\in\mathcal{H}^{\sigma}$, the map $\Phi$ is a contraction on
@@
for $T=T(A,v_{0})$ small enough, that is $T<CA^{-\frac{6(p-2)}{(p-8)}}$ for a constant $C>0$. For the Lipschitz continuity, consider two solutions $u$ and $v$ with respective initial data $u_{0}$ and $v_{0}$ in $H^{\sigma}$. We have
$\|u(t)-v(t)\|_{H^{\sigma}}$ $\lesssim\|e^{-it\mathcal{H}}(u_{0}-v_{0})\|_{H^{\sigma}}+\int_{0}^{t}\|e^{-i(t-s)\mathcal{H}}(|u(s)|^{p-2}u(s)-|v(s)|^{p-2}v(s))\|_{H^{\sigma}}\mathrm{d}s$ (3.29)
$\lesssim\|u_{0}-v_{0}\|_{H^{\sigma}}+\int_{0}^{t}\||u(s)|^{p-2}u(s)-|v(s)|^{p-2}v(s)\|_{H^{\sigma}}\mathrm{d}s$ (3.30)
$\lesssim\|u_{0}-v_{0}\|_{H^{\sigma}}+\int_{0}^{t}(1+\|u(s)\|_{H^{\sigma}}^{p-2}+\|v(s)\|_{H^{\sigma}}^{p-2})\|u(s)-v(s)\|_{H^{\sigma}}\mathrm{d}s$ (3.31)
hence Grönwall’s Lemma gives
$\|u(t)-v(t)\|_{H^{\sigma}}\leq\|u_{0}-v_{0}\|_{H^{\sigma}}e^{\int_{0}^{t}(1+\|u(s)\|_{H^{\sigma}}^{p-2}+\|v(s)\|_{H^{\sigma}}^{p-2})\mathrm{d}s}$ (3.32)
which completes the proof using that $\|u\|_{L^{\infty}([0,T],H^{\sigma})}\lesssim\|u_{0}\|_{H^{\sigma}}$ with the local well-posedness. To deal with $p\geq 8$, one needs to control higher integrability in time. Corollary 3.3 gives
@@
for any $p>8$ and $\frac{1}{q}=\frac{1-\theta}{2}+\frac{\theta}{6}$ with $\theta=\frac{6}{p-2}$. Similar computations as before give
$\|v(t)\|_{L^{p-2}([0,T],W^{\frac{1}{q}+\varepsilon,q})}\lesssim\|u_{0}\|_{H^{\frac{1-\kappa}{6}+\frac{1}{q}+\varepsilon}}+\|v\|_{L^{\infty}([0,T],H^{\frac{1-\kappa}{6}+\frac{1}{q}+\varepsilon})}\|v\|_{L^{p-2}([0,T],L^{\infty})}^{p-2}$ (3.34)
@@
With the relation between $q$ and $p$, we have
$\sigma_{\kappa}(p)=\frac{1-\kappa}{6}+\frac{1}{2}-\frac{2}{p-2}$ (3.36)
when $p>8$.
@@
$u(t)=e^{-it\Delta}u_{0}+i\int_{0}^{t}e^{-i(t-s)\Delta}Vu(s)\mathrm{d}s+i\lambda\int_{0}^{t}e^{-i(t-s)\Delta}|u(s)|^{p-2}u(s)\mathrm{d}s$ (3.37)
can not be used to deal with low regularity initial data due to the singular product $Vu$. This prevents the use of methods based on explicit space-time Fourier transform as done by Bourgain in [5] to prove local well-posedness since one has to deal with the Hamiltonian $\mathcal{H}$ and its Schrödinger group $e^{it\mathcal{H}}$ which is not explicit. Nevertheless our proof of local well-posedness does not rely too much on the form on the non-linearity and one can consider the truncated equation
$i\partial_{t}u_{N}=\mathcal{H}u_{N}+\lambda\Pi_{N}(|u_{N}|^{p-2}u_{N})$ (3.38)
with $u_{N}(0)=\Pi_{N}u_{0}$. Here $\Pi_{N}$ denotes the spectral projector associated to $\mathcal{H}$ on the space $E_{N}=\text{Vect}(e_{1},\ldots,e_{N})$ hence equation (3.38) is a finite dimensional system. Since $\Gamma$ is an isomorphism, the space $F_{N}:=\Gamma E_{N}$ is also of finite dimension. We have the following proposition.
@@
Let $V\in\mathcal{C}^{-1+\kappa}$ with $\kappa\in(0,1)$ and $p<8$. For any $\sigma\in(\sigma_{\kappa},\frac{1}{2})$, there exists a unique solution $u_{N}\in C([0,T],H^{\sigma})$ to equation (3.38) for $u_{0}\in H^{\sigma}$ up to time $T$ of order $\|u_{0}\|_{H^{\sigma}}^{-\frac{6(p-2)}{p-8}}$ uniform with respect to $N\geq 1$ and we have
$\sup_{t\in[0,T]}\|u_{N}(t)\|_{H^{\sigma}}\leq 2\|u_{0}\|_{H^{\sigma}}.$ (3.39)
Moreover, we have
$\sup_{t\in[0,T]}\|u_{N}(t)-u(t)\|_{H^{\sigma^{\prime}}}\lesssim N^{\sigma^{\prime}-\sigma}(1+\|u_{0}\|_{H^{\sigma}}^{p-2})$ (3.40)
@@
for any $\sigma^{\prime}<\sigma$ where $u$ is the solution to the untruncated equation (1.1) with initial data $u_{0}$.
Proof : As for the untruncated equation, we consider the new variable $v_{N}=\Gamma^{-1}u_{N}$ which satisfies the mild formulation
$v_{N}(t)=e^{-it\mathcal{H}^{\sharp}}\Pi_{N}v_{0}+i\lambda\int_{0}^{t}e^{-i(t-s)\mathcal{H}^{\sharp}}\Gamma^{-1}\Pi_{N}\big{(}|\Gamma v_{N}(s)|^{p-2}\Gamma_{N}v(s)\big{)}\mathrm{d}s$ (3.41)
with initial data $v_{N}(0)=\Pi_{N}\Gamma^{-1}u_{0}$. Since $\Pi_{N}$ is continuous from $\mathcal{D}^{\sigma}$ to itself for any $\sigma\in\mathbb{R}$ uniformly with respect to $N\geq 1$, it is continuous from $H^{\sigma}$ to itself for $|\sigma|\leq 1+\kappa$. Thus the fixed point argument applies also to the truncated equation with a solution up to a time $T$ uniform with respect to $N\geq 1$ of order $\|u_{0}\|_{H^{\sigma}}^{-\frac{6(p-2)}{p-8}}$ with
$\sup_{t\in[0,T]}\|u_{N}(t)\|_{H^{\sigma}}\leq 2\|u_{0}\|_{H^{\sigma}}.$ (3.42)
For the convergence, we consider the difference $w_{N}:=u-u_{N}$ which satisfies the equation
$i\partial_{t}w_{N}=\mathcal{H}w_{N}+\lambda|u|^{p-2}u-\lambda\Pi_{N}(|u_{N}|^{p-2}u_{N})$ (3.43)
with initial condition $w_{N}(0)=(\text{Id}-\Pi_{N})u_{0}$. The mild formulation yields
[…]
$\lesssim N^{\sigma-\sigma^{\prime}}\|u_{0}\|_{H^{\sigma}},+\int_{0}^{t}\big{\|}|u|^{p-2}u-\Pi_{N}(|u|^{p-2}u)\big{\|}_{H^{\sigma^{\prime}}}\mathrm{d}s+\int_{0}^{t}\big{\|}\Pi_{N}(|u|^{p-2}u-|u_{N}|^{p-2}u_{N})\big{\|}_{H^{\sigma^{\prime}}}\mathrm{d}s$
$\lesssim N^{\sigma-\sigma^{\prime}}\|u_{0}\|_{H^{\sigma}},+\int_{0}^{t}N^{\sigma-\sigma^{\prime}}\big{\|}|u|^{p-2}u\big{\|}_{H^{\sigma}}\mathrm{d}s+\int_{0}^{t}\big{\|}|u|^{p-2}u-|u_{N}|^{p-2}u_{N}\big{\|}_{H^{\sigma^{\prime}}}\mathrm{d}s$
$\lesssim N^{\sigma-\sigma^{\prime}}\|u_{0}\|_{H^{\sigma}},+\int_{0}^{t}N^{\sigma-\sigma^{\prime}}\|u\|_{L^{\infty}}^{p-2}\|u\|_{H^{\sigma}}\mathrm{d}s+\int_{0}^{t}(1+\|u\|_{L^{\infty}}^{p-2}+\|u_{N}\|_{L^{\infty}}^{p-2})\|w_{N}\|_{H^{\sigma^{\prime}}}\mathrm{d}s$
$\lesssim N^{\sigma-\sigma^{\prime}}\|u_{0}\|_{H^{\sigma}},+N^{\sigma-\sigma^{\prime}}\|u\|_{\mathcal{S}_{T}}^{p-1}+T^{\frac{8-p}{6}}(1+\|u\|_{\mathcal{S}_{T}}^{p-2}+\|u_{N}\|_{\mathcal{S}_{T}}^{p-2})\sup_{t\in[0,T]}\|w_{N}(t)\|_{H^{\sigma^{\prime}}}$
for any $\sigma^{\prime}<\sigma$. Since we have
$\|u_{N}\|_{\mathcal{S}_{T}}+\|u\|_{\mathcal{S}_{T}}\lesssim\|u_{0}\|_{H^{\sigma}},$ (3.44)
we get
$\sup_{t\in[0,T]}\|w_{N}(t)\|_{H^{\sigma}}\lesssim N^{\sigma-\sigma^{\prime}}(1+\|u_{0}\|_{H^{\sigma}}^{p-2})$ (3.45)
for $T$ of order $\|u_{0}\|_{H^{\sigma}}^{-\frac{6(p-2)}{p-8}}$.
@@
$\mathcal{E}(u)=\int_{\mathbb{T}}|\partial_{x}u(x)|^{2}\mathrm{d}x+\int_{\mathbb{T}}|u(x)|^{2}V(\mathrm{d}x)+\lambda\int_{\mathbb{T}}|u(x)|^{p}\mathrm{d}x$ (4.2)
@@
$\nu(\mathrm{d}u)=\frac{1}{Z}e^{(1-\lambda_{1})\int_{\mathbb{T}}|u(x)|^{2}\mathrm{d}x-\lambda\int_{\mathbb{T}}|u(x)|^{p}\mathrm{d}x}\mu(\mathrm{d}u)$ (4.5)
with the normalisation constant $Z=\|e^{(1-\lambda_{1})\int_{\mathbb{T}}|u(x)|^{2}\mathrm{d}x-\lambda\int_{\mathbb{T}}|u(x)|^{p}\mathrm{d}x}\|_{L^{1}(\mu)}$. The measure $\mu$ corresponds to the law of the Gaussian Free Field (GFF) associated to $\mathcal{H}-\lambda_{1}+1$, that is the random serie
@@
Since the measure is supported on continuous functions, the potential energy $\|u\|_{L^{p}(\mathbb{T})}^{p}$ is finite for $\mu$-almost all functions and the definition of the measure
$\nu(\mathrm{d}u)=\frac{1}{Z}\exp\Big{(}(1-\lambda_{1})\int_{\mathbb{T}}|u(x)|^{2}\mathrm{d}x-\lambda\int_{\mathbb{T}}|u(x)|^{p}\mathrm{d}x\Big{)}\mu(\mathrm{d}u)$ (4.18)
amounts to proving $\exp\Big{(}(1-\lambda_{1})\int_{\mathbb{T}}|u(x)|^{2}\mathrm{d}x-\lambda\int_{\mathbb{T}}|u(x)|^{p}\mathrm{d}x\Big{)}\in L^{1}(\mu)$. In the defocusing case $\lambda>0$, this follows from the fact that $\mu$ is a probability measure since the exponential is bounded for $p>2$ using the injection $L^{p}\hookrightarrow L^{2}$. Indeed, one has
$\exp\Big{(}(1-\lambda_{1})\int_{\mathbb{T}}|u(x)|^{2}\mathrm{d}x-\lambda\int_{\mathbb{T}}|u(x)|^{p}\mathrm{d}x\Big{)}$ $\leq\exp\Big{(}(1-\lambda_{1})\int_{\mathbb{T}}|u(x)|^{2}\mathrm{d}x-\lambda c_{p}\big{(}\int_{\mathbb{T}}|u(x)|^{2}\mathrm{d}x\big{)}^{\frac{p}{2}}\Big{)}$ (4.19)
$\leq\sup_{r>0}e^{(1-\lambda_{1})r^{2}-\lambda c_{p}r^{p}}$ (4.20)

### v2

@@ Abstract
$\sigma_{\kappa}(m)=\frac{1}{3}-\frac{\kappa}{6}$ (3.13)
when $m\leq 8$ and
$\sigma_{\kappa}(m)=\frac{4-\kappa}{6}-\frac{2}{m-2}$ (3.14)
when $m>8$.
@@
Let $V\in\mathcal{C}^{-1+\kappa}$ and $m\geq 2$ an integer. For any $\sigma\in(\sigma_{\kappa}(m),1+\kappa]$, there exists a unique solution $u\in C([0,T],H^{\sigma})$ to equation (1.1) for $u_{0}\in H^{\sigma}$ up to a time $T$ of order $\|u_{0}\|_{H^{\sigma}}^{-\frac{6(m-2)}{m-8}}$ and we have
$\|u\|_{L^{6}([0,T],L^{\infty})}+\sup_{t\in[0,T]}\|u(t)\|_{H^{\sigma}}\leq 2\|u_{0}\|_{H^{\sigma}}.$ (3.15)
Moreover, the solution is locally Lipschitz with respect to the initial data from $H^{\sigma}$ to $C([0,T],H^{\sigma})\cap L^{6}([0,T],L^{\infty})$.
@@
$i\partial_{t}v=\mathcal{H}^{\sharp}v+\lambda\Gamma^{-1}(|\Gamma v|^{m-2}\Gamma v)$ (3.16)
with initial data $v(0)=\Gamma^{-1}u_{0}\in\Gamma^{-1}H^{\sigma}=H^{\sigma}$ since $\sigma\leq 1+\kappa$. We prove existence and uniqueness of solutions for this equation which gives a unique solution to the initial equation using that $\Gamma$ is continuous from $H^{\sigma}$ to itself. Note that this is direct for $\sigma\in(\frac{1}{2},1+\kappa]$ using the continuity of $\Gamma$ and its inverse and that $H^{\sigma}$ is an algebra. For the rest of the proof, we consider $\sigma\leq\frac{1}{2}$. For $T>0$, let $\mathcal{S}_{T}$ be the solution space
@@
for $2\varepsilon\in(0,\sigma-\sigma_{\kappa}]$. We prove that the map
$\Phi(v)(t):=e^{-it\mathcal{H}^{\sharp}}v_{0}-i\lambda\int_{0}^{t}e^{-i(t-s)\mathcal{H}^{\sharp}}\Gamma^{-1}\big{(}|\Gamma v(s)|^{m-2}\Gamma v(s)\big{)}\mathrm{d}s$ (3.19)
@@
$\|\Phi(v)(t)\|_{H^{\sigma}}$ $\lesssim\|v_{0}\|_{H^{\sigma}}+\int_{0}^{t}\||\Gamma v(s)|^{m-2}\Gamma v(s)\|_{H^{\sigma}}\mathrm{d}s$
$\lesssim\|v_{0}\|_{H^{\sigma}}+\int_{0}^{t}\|v(s)\|_{H^{\sigma}}\|v(s)\|_{L^{\infty}}^{m-2}\mathrm{d}s$
$\lesssim\|v_{0}\|_{H^{\sigma}}+\|v\|_{L^{\infty}([0,t],H^{\sigma})}T^{1-\frac{m-2}{6}}\|v\|_{L^{6}([0,t],L^{\infty})}^{m-2}$
using that $\Gamma$ is continuous from $H^{\sigma}$ to itself, from $L^{\infty}$ to itself and Hölder inequality for $m<8$ with $t\in[0,T]$. Since $\sigma\leq\frac{1}{2}$, the space $H^{\sigma}$ is not an algebra and one needs to control the $L^{\infty}$ norm, this is where Strichartz inequalities are crucial with the embedding
@@
$\sup_{t\in[0,T]}\|\Phi(v)(t)\|_{H^{\sigma}}\lesssim\|v_{0}\|_{H^{\sigma}}+T^{\frac{8-m}{6}}\|v\|_{\mathcal{S}_{T}}^{m-1}.$ (3.21)
Using that the spaces $\mathcal{D}^{\delta}$ and $H^{\delta}$ are equivalent for $|\delta|\leq 1+\kappa$ and the continuity results on $\Gamma$, Theorem 3.1 gives
@@
$\|\Phi(v)\|_{L^{6}([0,t],W^{\frac{1}{6}+\varepsilon,6})}$ $\lesssim\|v_{0}\|_{H^{\frac{1-\kappa}{6}+2\varepsilon+\frac{1}{6}}}+\int_{0}^{t}\|v(s)\|_{H^{\frac{1-\kappa}{6}+2\varepsilon+\frac{1}{6}}}\|v(s)\|_{L^{\infty}}^{m-2}\mathrm{d}s$ (3.23)
$\lesssim\|v_{0}\|_{H^{\frac{1-\kappa}{6}+2\varepsilon+\frac{1}{6}}}+\|v\|_{L^{\infty}([0,t],H^{\frac{1-\kappa}{6}+2\varepsilon+\frac{1}{6}})}T^{1-\frac{m-2}{6}}\|v\|_{L^{6}([0,t],W^{\frac{1}{6}+\varepsilon,6})}^{m-2}$ (3.24)
$\lesssim\|v_{0}\|_{H^{\sigma}}+T^{\frac{8-m}{6}}\|v\|_{\mathcal{S}_{T}}^{m-1}$ (3.25)
since $\sigma\geq\sigma_{\kappa}+2\varepsilon$. This gives
$\|\Phi(v)\|_{\mathcal{S}_{T}}\lesssim\|v_{0}\|_{H^{\sigma}}+T^{\frac{8-m}{6}}\|v\|_{\mathcal{S}_{T}}^{m-1}$ (3.26)
hence for any $A\geq 0$ large enough depending on the initial condition, the ball centered in 00 of size $A$ is stable by $\Phi$ for $T>0$ small enough. Similar computations give
$\|\Phi(v)-\Phi(v^{\prime})\|_{\mathcal{S}_{T}}\lesssim T^{\frac{8-m}{6}}(1+\|v\|_{\mathcal{S}_{T}}^{m-2}+\|v^{\prime}\|_{\mathcal{S}_{T}}^{m-2})\|v-v^{\prime}\|_{\mathcal{S}_{T}}$ (3.27)
for $v,v^{\prime}\in\mathcal{S}_{T}$ with $v(0)=v^{\prime}(0)$. This gives that for $v_{0}\in\mathcal{H}^{\sigma}$ and $A=A(v_{0})$ large enough, the map $\Phi$ is a contraction on
@@
for $T=T(A,v_{0})$ small enough, that is $T<CA^{-\frac{6(m-2)}{(m-8)}}$ for a constant $C>0$. This also gives that the flow map $u_{0}\mapsto u$ is Lipschitz continuous from bounded subsets of $H^{\sigma}$ to $C([0,T],H^{\sigma})\cap L^{6}([0,T],L^{\infty})$.
To deal with $m\geq 8$, one needs to control higher integrability in time. Corollary 3.3 gives
@@
for any $m>8$ and $\frac{1}{q}=\frac{1-\theta}{2}+\frac{\theta}{6}$ with $\theta=\frac{6}{m-2}$. Similar computations as before give
$\|v(t)\|_{L^{m-2}([0,T],W^{\frac{1}{q}+\varepsilon,q})}\lesssim\|u_{0}\|_{H^{\frac{1-\kappa}{6}+\frac{1}{q}+\varepsilon}}+\|v\|_{L^{\infty}([0,T],H^{\frac{1-\kappa}{6}+\frac{1}{q}+\varepsilon})}\|v\|_{L^{m-2}([0,T],L^{\infty})}^{m-2}$ (3.29)
@@
With the relation between $q$ and $m$, we have
$\sigma_{\kappa}(m)=\frac{1-\kappa}{6}+\frac{1}{2}-\frac{2}{m-2}$ (3.31)
when $m>8$.
@@
$u(t)=e^{-it\Delta}u_{0}+i\int_{0}^{t}e^{-i(t-s)\Delta}Vu(s)\mathrm{d}s-i\lambda\int_{0}^{t}e^{-i(t-s)\Delta}|u(s)|^{m-2}u(s)\mathrm{d}s$ (3.32)
can not be used to deal with low regularity initial data due to the singular product $Vu$. This prevents the use of methods based on explicit space-time Fourier transform as done by Bourgain in [5] to prove local well-posedness since one has to deal with the Hamiltonian $\mathcal{H}$ and its Schrödinger group $e^{it\mathcal{H}}$ which are not explicit. Nevertheless our proof of local well-posedness does not rely too much on the form of the non-linearity and one can consider the truncated equation
$i\partial_{t}u_{N}=\mathcal{H}u_{N}+\lambda\Pi_{N}(|u_{N}|^{m-2}u_{N})$ (3.33)
with $u_{N}(0)=\Pi_{N}u_{0}$. Here $\Pi_{N}$ denotes the spectral projector associated to $\mathcal{H}$ on the space $E_{N}=\text{Vect}(e_{1},\ldots,e_{N})$ hence equation (3.33) is a finite dimensional system. Since $\Gamma$ is an isomorphism, the space $F_{N}:=\Gamma E_{N}$ is also of finite dimension. We have the following proposition which states that the Galerkin approximation converges to the maximal solution on any time interval.
@@
Let $V\in\mathcal{C}^{-1+\kappa}$ with $\kappa\in(0,1)$ and $m<8$ an integer. For any $\sigma\in(\sigma_{\kappa},\frac{1}{2})$, let $u$ be the unique maximal solution $u\in C([0,T^{*}(u_{0})),H^{\sigma})$ to equation (1.1) for $u_{0}\in H^{\sigma}$. Then for any $T<T^{*}(u_{0})$ and $\sigma^{\prime}<\sigma$, there exists a constant $C>0$ independent of $N$ such that
$\|u-u_{N}\|_{L^{6}([0,T],L^{\infty})}+\sup_{t\leq T}\|u(t)-u_{N}(t)\|_{H^{\sigma^{\prime}}}\leq CN^{\sigma^{\prime}-\sigma}$ (3.34)
@@
for $N$ large enough and $u_{N}$ the solution to the truncated equation (3.33) with initial data $\Pi_{N}u_{0}$.
Proof : The same fixed point argument for the truncated equation (3.33) gives an existence up to a time $T=CA^{-c}>0$ for $\|u_{N}(0)\|_{H^{\sigma}}\leq A$ with uniform bound
$\|u_{N}\|_{L^{6}([0,T],L^{\infty})}+\sup_{t\in[0,T]}\|u_{N}(t)\|_{H^{\sigma}}\leq 2A.$ (3.35)
Similar computations also gives
$\|u-u_{N}\|_{L^{6}([0,T],L^{\infty})}+\sup_{t\in[0,T]}\|u(t)-u_{N}(t)\|_{H^{\sigma}}\leq C_{A}\|u(0)-u_{N}(0)\|_{H^{\sigma}}$ (3.36)
for a constant $C_{A}>0$ since the solution is locally Lipschitz continuous with respect to the initial data.
Now consider the global solution $u$ to equation (1.1) defined on $[0,T^{*}(u_{0}))$ with $T^{*}(u_{0})\in(0,\infty]$ for $u_{0}\in H^{\sigma}$ and $\sigma^{\prime}<\sigma$. Consider also the solution $u_{N}$ to the truncated equation (3.33) with initial data $u_{N}(0)=\Pi_{N}u_{0}$. Let $T<T^{*}(u_{0})$ and
$B:=b\sup_{t\in[0,T]}\|u(t)\|_{H^{\sigma}}+1$ (3.37)
with $b=\max(1,\|\Pi_{N}\|_{H^{\sigma}\to H^{\sigma}})>0$. Then we have $\|u_{0}\|_{H^{\sigma}}\leq B$ and $\|\Pi_{N}u_{0}\|_{H^{\sigma}}\leq B$ thus
$\|u-u_{N}\|_{L^{6}([0,\tau],L^{\infty})}+\sup_{t\in[0,\tau]}\|u(t)-u_{N}(t)\|_{H^{\sigma}}\leq C_{B}\|u_{0}-\Pi_{N}u_{0}\|_{H^{\sigma}}$ (3.38)
for $\tau=CB^{-c}>0$ and $C_{B}>0$. Since $\sigma^{\prime}<\sigma$, we get
$\sup_{t\in[0,\tau]}\|u(t)-u_{N}(t)\|_{H^{\sigma^{\prime}}}\leq C_{B}N^{\sigma^{\prime}-\sigma}\|u_{0}\|_{H^{\sigma}}$ (3.39)
thus
$\|u_{N}(\tau)\|_{H^{\sigma^{\prime}}}\leq\|u(\tau)\|_{H^{\sigma^{\prime}}}+C_{B}N^{\sigma^{\prime}-\sigma}\|u_{0}\|_{H^{\sigma}}.$ (3.40)
For $N$ large enough, we get
$\|u_{N}(\tau)\|_{H^{\sigma^{\prime}}}\leq\|u(\tau)\|_{H^{\sigma}}+1\leq B.$ (3.41)
Thus we have $\|u(\tau)\|_{H^{\sigma^{\prime}}}\leq B$ as well as $\|u_{N}(\tau)\|_{H^{\sigma^{\prime}}}\leq B$ and the same argument gives
$\|u-u_{N}\|_{L^{6}([\tau,2\tau],L^{\infty})}+\sup_{t\in[\tau,2\tau]}\|u(t)-u_{N}(t)\|_{H^{\sigma^{\prime}}}\leq C_{B}\|u(\tau)-u_{N}(\tau)\|_{H^{\sigma^{\prime}}}$ (3.42)
thus
$\sup_{t\in[0,2\tau]}\|u(t)-u_{N}(t)\|_{H^{\sigma^{\prime}}}\leq(1+C_{B})C_{B}N^{\sigma^{\prime}-\sigma}\|u_{0}\|_{H^{\sigma}}.$ (3.43)
Again, taking $N$ large enough yields
$\|u_{N}(2\tau)\|_{H^{\sigma^{\prime}}}\leq\|u(2\tau)\|_{H^{\sigma^{\prime}}}+1\leq B$ (3.44)
thus a finite number of repition of this argument yields
$\|u-u_{N}\|_{L^{6}([0,T],L^{\infty})}+\sup_{t\in[0,T]}\|u(t)-u_{N}(t)\|_{H^{\sigma^{\prime}}}\leq CN^{\sigma^{\prime}-\sigma}\|u_{0}\|_{H^{\sigma}}$ (3.45)
for a constant $C>0$.
@@
$\mathcal{E}(u)=\int_{\mathbb{T}}|\partial_{x}u(x)|^{2}\mathrm{d}x+\int_{\mathbb{T}}|u(x)|^{2}V(\mathrm{d}x)+\lambda\int_{\mathbb{T}}|u(x)|^{m}\mathrm{d}x$ (4.2)
@@
$\nu(\mathrm{d}u)=\frac{1}{Z}e^{(1-\lambda_{1})\int_{\mathbb{T}}|u(x)|^{2}\mathrm{d}x-\lambda\int_{\mathbb{T}}|u(x)|^{m}\mathrm{d}x}\mu(\mathrm{d}u)$ (4.5)
with the normalisation constant $Z=\|e^{(1-\lambda_{1})\int_{\mathbb{T}}|u(x)|^{2}\mathrm{d}x-\lambda\int_{\mathbb{T}}|u(x)|^{m}\mathrm{d}x}\|_{L^{1}(\mu)}$. The measure $\mu$ corresponds to the law of the Gaussian Free Field (GFF) associated to $\mathcal{H}-\lambda_{1}+1$, that is the random serie
@@
Since the measure is supported on continuous functions, the potential energy $\|u\|_{L^{m}(\mathbb{T})}^{m}$ is finite for $\mu$-almost all functions and the definition of the measure
$\nu(\mathrm{d}u)=\frac{1}{Z}\exp\Big{(}(1-\lambda_{1})\int_{\mathbb{T}}|u(x)|^{2}\mathrm{d}x-\lambda\int_{\mathbb{T}}|u(x)|^{m}\mathrm{d}x\Big{)}\mu(\mathrm{d}u)$ (4.18)
amounts to proving $\exp\Big{(}(1-\lambda_{1})\int_{\mathbb{T}}|u(x)|^{2}\mathrm{d}x-\lambda\int_{\mathbb{T}}|u(x)|^{m}\mathrm{d}x\Big{)}\in L^{1}(\mu)$. In the defocusing case $\lambda>0$, this follows from the fact that $\mu$ is a probability measure since the exponential is bounded for $m>2$ using the injection $L^{m}\hookrightarrow L^{2}$. Indeed, one has
$\exp\Big{(}(1-\lambda_{1})\int_{\mathbb{T}}|u(x)|^{2}\mathrm{d}x-\lambda\int_{\mathbb{T}}|u(x)|^{m}\mathrm{d}x\Big{)}$ $\leq\exp\Big{(}(1-\lambda_{1})\int_{\mathbb{T}}|u(x)|^{2}\mathrm{d}x-\lambda c_{m}\big{(}\int_{\mathbb{T}}|u(x)|^{2}\mathrm{d}x\big{)}^{\frac{m}{2}}\Big{)}$ (4.19)
$\leq\sup_{r>0}e^{(1-\lambda_{1})r^{2}-\lambda c_{m}r^{m}}$ (4.20)